A spherical vessel used for deep-sea exploration has a radius of 1.52 m and a mass of 1.16 104 kg. To dive, the vessel takes on mass in the form of sea water. Determine the amount of mass that the vessel must take on if it is to descend at a constant speed of 1.10 m/s, when the resistive force on it is 1099 N in the upward direction. The density of seawater is 1.03 103 kg/m3. kg
Answer #1

net mass of submarine

Ms = density of water * volume of submarine - Fr / g

Ms = 1030* (4 / 3)* pi 1.52^3 - 1099 / 9.8

Ms = 1.503* 10^4 kg

now

mass that submarine can take

Ms = m + ms

1.503* 10^4 = m + 1.16* 10^4

m = 3430 kg

========